OpenStudy (brucebaner):

The altitude of a plane as it begins to descend is shown by the linear equation a = 22,000 – 800m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 12? ---------

OpenStudy (brucebaner):

@KamiBug

OpenStudy (brucebaner):

I think a= -9578 , but I'm not sure

OpenStudy (brucebaner):

@confluxepic

OpenStudy (brucebaner):

@KamiBug

OpenStudy (kamibug):

a = 22,000 – 800(12) a = 22,000 - 9600 a = ?

OpenStudy (brucebaner):

or, 12 x -800 = 9600 so a = 22000 - 9600 = 12400

OpenStudy (brucebaner):

so em i right @KamiBug

OpenStudy (kamibug):

The answer is 12,400, yes. :)